OpenStudy (anonymous):

if f(x)= square root of x+12 and g(x)=2 square root of x, what if the value of (f-g)(144)

ganeshie8 (ganeshie8):

\(\large f(x) = \sqrt{x+12}\) \(\large g(x) = 2\sqrt{x}\)

ganeshie8 (ganeshie8):

like that ?

OpenStudy (anonymous):

Yes!

ganeshie8 (ganeshie8):

or may be like this : \(\large f(x) = \sqrt{x} + 12\) \(\large g(x) = 2\sqrt{x}\)

OpenStudy (anonymous):

First one.

OpenStudy (anonymous):

wait, no. The second.

ganeshie8 (ganeshie8):

haha yes the second one makes more sense

ganeshie8 (ganeshie8):

use below : \(\large (f-g)(144) = f(144) - g(144)\)

ganeshie8 (ganeshie8):

evaluate f(x) at x = 144 evaluate g(x) at x = 144 and subtract

OpenStudy (anonymous):

So, 0.

ganeshie8 (ganeshie8):

Excellent !

OpenStudy (anonymous):

Oh great thanks. Haha. 6 more questions to go..

ganeshie8 (ganeshie8):

np :) good luck with your homework !